frank and sam have ages that are consecutive odd integers. the product of their ages is 899. which equation could be used to find franks age if he is the younger one
Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

frank and sam have ages that are consecutive odd integers.
the product of their ages is 899.
which equation could be used to find franks age if he is the younger one
:
x = Frank's age
x(x+2) = 899
x^2 + 2x - 899 = 0
Factors to
(x+31)(x-29) = 0
x = 29 is franks age

Answer by Alan3354(69443)

frank and sam have ages that are consecutive odd integers. the product of their ages is 899. which equation could be used to find franks age if he is the younger one
------------
(n-1)*(n+1) = 899
n^2 - 1 = 899
n = 900
n = 30
--> 29 & 31